Assume that a sample is used to estimate a population proportion p. Find the 80% confidence interval for a sample of size 227 with 112 successes. Enter your answer as a tri-linear inequality using decimals (not percents) accurate to three decimal places.

____< p < ____
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 112 / 227 = 0.493

1 - \hat p = 1n - 0.493 = 0.507

Z\alpha/2 = 1.282

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.282 * (\sqrt((0.493 * 0.507) / 227)

Margin of error = E = 0.043

A 80% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.493 - 0.043 < p < 0.493 + 0.043

0.450 < p < 0.536